OpenStudy (anonymous):

Integral of (lnx) / x ^ 2 from 1 to 2

OpenStudy (anonymous):

I couldn't find my original post so I had to repost it :/

OpenStudy (anonymous):

integrate by parts

OpenStudy (anonymous):

I know that but how!?

OpenStudy (anonymous):

choose f and dg so that the integral df g will solve easily

OpenStudy (anonymous):

let f be lnx and dg be 1/x^2

OpenStudy (anonymous):

df = dx/x and g= -1/x

OpenStudy (anonymous):

\[\int\limits_{}^{}fdg = fg - \int\limits_{}^{}gdf\]

OpenStudy (anonymous):

But I end up taking the integral of a natural log... don't i?

OpenStudy (anonymous):

no, you end up taking the derivative of it

OpenStudy (anonymous):

Oh you used the opposite for f and dg... ok that works